# Color: A Framework for Applying Graph Coloring to Subgraph Cardinality Estimation

### v1

@@ Abstract.
* These authors contributed equally to the work.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Special Interest Group on the Management of Data; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 4.2. Special Case: Stable Colorings
This theorem states that stable colorings are a perfect statistic for cardinality estimation of acyclic query graphs. However, we cannot use them in practice, because the number of stable colors needed to represent real graphs is close to the number of vertices in the data graph (Morris et al., 2021; Kayali and Suciu, 2022). However, the theorem demonstrates that, as colorings approach stability, the estimate converges to the true subgraph count.
@@ 7.1. Partial Aggregation
Naively, the runtime of inference on the lifted graph would be exponential in the size of the query graph, making it intractable for even moderately sized query graphs. This is because the lifted graph is complete, or nearly complete, so the size of $Hom(Q,F)$ is approximately $|C|^{|Q|}$. Fortunately, we can rephrase the sum in Eq. 3 to drastically reduce this runtime via aggregate push-down. To do this, we express the set of lifted graph matches, $\text{hom}(Q,F)$, as $(c_{1},\ldots,c_{|Q|})\in C^{|Q|}$. We then use the definition of $W$ from 6 and, as before, we assume a topological ordering on the edges, $e_{1},\ldots,e_{|E|}$, and vertices, $v_{1},\ldots,v_{|Q|}$.
@@
At this point, we can identify this problem as a Functional Aggregate Query as described in (Khamis et al., 2016), and we can apply the techniques there to solve it efficiently. [^2] Note, this is closely related to the variable elimination algorithm for probabilistic graphical models as well as tensor contraction algorithms. As an example, suppose that the query graph $Q$ is a line graph $v_{1}\rightarrow v_{2}\rightarrow v_{3}$. The naive expression would be the following $O(|C|^{3})$ expression,
@@
More generally, we can apply this strategy by choosing a variable order and at each step summing out the next variable in the order. The efficacy of this strategy depends on the variable order, and, in particular, it depends on the maximum number of variables present in any intermediate product. If an intermediate product involves $k$ variables, then we need to compute a relation of size $|C|^{k}$ which dominates the runtime. We defer the details of the proof to the technical report (tec, 2024), but this intuition can be formalized as follows using the theory of tree decompositions and treewidth,
@@
Of course, this relies on finding a good ordering of the query vertices, and finding the optimal one is naively NP-Hard with respect to the size of $Q$. Fortunately, there are very effective heuristics for identifying good tree decompositions, and we apply these in our system, using the min-fill heuristic (Rollon and Larrosa, 2011). Further, in order to accommodate the sampling techniques that we discuss next, we restrict these tree decompositions to path decompositions and get runtime results relative to the pathwidth.
7.2. Sampling Techniques
In real world systems, cardinality estimation needs to be an extremely fast and consistent process because it is an overhead incurred by every query. While partial aggregation significantly speeds up inference for simple query graphs with low treewidth, larger and denser query graphs may still pose a problem under these requirements. To avoid this, we integrate a sampling procedure into the inference process which ensures a linear inference time w.r.t the size of the query (see Fig. 13).
At a high level, this sampling procedure is similar to a weighted version of the WanderJoin algorithm from (Li et al., 2019). We apply a Thompson-Horowitz estimator to randomly chosen paths within the lifted graph. However, we adjust the method in two important ways: 1) we incorporate the sampling into the aggregation framework from Sec. 7.1 2) we apply importance sampling to account for the fact that different paths within the lifted graphs contribute more or less to the cardinality estimate.
Sampling During Aggregation
At each step of our algorithm, we process a single vertex of the query graph and materialize an intermediate result consisting of partial colorings and weights associated with them. After this materialization, we apply sampling in order to reduce the amount of partial colorings that we extend in the next step. By doing this at each step, we can maintain a constant number of partial colorings at all times and ensure a linear runtime w.r.t. the size of the query graph. Further, because we can still apply aggregation where possible, we reduce the amount of sampling required to maintain a small number of partial colorings.
Importance Sampling
This is a classical technique for approximating the value of an integral, and we adapt it here by noting that our summation in 8 is simply a discrete integral over a product. The core idea is to sample points of the integrand which contribute more heavily to the result with higher probability in order to reduce the variance. Because determining the contribution of a partial coloring to the final result is inherently challenging, we instead approximate this contribution via its associated weight. This reflects an assumption that partial colorings with high weight are likely to disproportionately contribute to the final sum. To keep our estimator unbiased, we apply Thompson-Horowitz estimation and simply multiply the weight of each sampled partial coloring by the inverse of its selection probability. Finally, we scale the total weight of the sampled colorings to make it equal to the weight of the partial colorings prior to sampling.
@@ 7.2. Sampling Techniques
human 0.67 0.00 0.22 0.63 0.00 0.00 0.00 0.00 0.00 0.00
aids 0.69 0.07 0.14 0.28 0.00 0.04 0.01 0.02 0.39 0.00
lubm80 0.83 0.17 0.67 0.67 0.00 0.00 0.00 0.00 0.00 0.00
yeast 1.00 0.97 0.97 0.11 0.00 0.63 0.60 0.63 0.88 0.00
dblp 1.00 0.99 0.94 0.15 0.00 0.14 0.14 0.70 0.85 0.00
youtube 0.99 0.93 1.00 0.22 0.00 0.10 0.05 0.63 0.78 0.00
eu2005 0.95 0.90 0.91 0.55 0.00 0.00 0.00 0.22 0.44 0.00
patents 0.98 0.88 0.98 0.08 0.00 0.13 0.13 0.67 0.79 0.00
@@ 7.3. Handling Updates
Updates pose a challenge to summary-based estimators because the statistics which they collect become stale over time as updates are applied to the database. Traditionally, summary based estimators simply recalculate the summary on a regular basis to accommodate updates (Haas, 1999). This approach leads to severe decreases in accuracy under even modest updates because the estimator is entirely ”blind” to them. On the other hand, recalculating the summary before each query is far too costly. In this section, we demonstrate how COLOR supports a middle ground approach that applies fast, basic updates to the lifted graph, allowing it to maximize the time between full rebuilds.
@@
Given an update $\theta_{E}=(v_{1},v_{2},A)$, the edge update function $\delta_{E}$ marginally increases $\tau$ for the combination of attributes and colors in the edge update. However, the edge update doesn’t directly contain the colors of $v_{1}$ and $v_{2}$ or the attributes of $v_{2}$. To retrieve the colors, $c_{1}$ and $c_{2}$, associated with $v_{1}$ and $v_{2}$, we look up their values in $\pi$ which we store compactly (and approximately) as a series of cuckoo filters. To calculate the attributes of $v_{2}$, we leverage statistics about the attribute distribution in $c_{2}$ and sample a set of attributes from that distribution.
@@ 8. Evaluation
    Never experiences estimation failure on any query across all workloads—unlike all but one comparison method.
@@
    Produced up to $\mathbf{10^{7}\times}$ tighter bounds when using max degree statistics, $\tau_{max}$, than BoundSketch while being many orders of magnitude faster.
@@
Datasets & Workload. We consider datasets from (Park et al., 2020a) and (Sun and Luo, 2020) for our analysis. These datasets come from a variety of domains. Those from (Sun and Luo, 2020) are undirected graphs whose queries are larger and vary in density. Those from (Park et al., 2020a) are directed graphs whose queries are smaller and vary in shape. We adapt undirected workloads to directed methods by including reverse edges in the data graph but not in the query graphs. Both of these sources include label predicates in their queries with the former using both edge and vertex labels and the latter using only vertex labels. Table 3 shows their different characteristics where $|\ell_{V}|$ and $|\ell_{E}|$ refer to the number of unique edge and vertex labels.
Comparison Methods. For comparison, we use a superset of the methods considered in (Park et al., 2020a) and additionally apply them to the larger, more complex datasets from (Sun and Luo, 2020). These methods include: 1) Correlated Sampling (CS) (Vengerov et al., 2015) 2) Characteristic Sets (CSet)(Neumann and Moerkotte, 2011) 3) Wander Join (WJ) (Li et al., 2019) 4) Alley (alley) and alleyTPI (Kim et al., 2021) 5) Join Sampling with Upper Bounds (JSUB), an adaptation of (Zhao et al., 2018) 6) Bound Sketch (BSK) (Cai et al., 2019) which corresponds to a hash-based coloring and using the max degree estimator. We further apply our partial aggregation optimization, and we call this improved version BSK++. 7) IMPR (Chen and Lui, 2018) and 8) SumRDF (Stefanoni et al., 2018) 9) We also include a traditional independence-based estimator (IndEst) corresponding to Def. 2.4. For the sampling based estimators, we apply the default sampling ratios from (Park et al., 2020a) and (Kim et al., 2021) (i.e. .03 for all methods except for Alley which uses .001). AlleyTPI uses a maximum pattern length (MAX_L) and a maximum number of stored label groups (NUM_GROUPS) when building its index, with default values of MAX_L=4 or MAX_L=5 depending on the dataset, and NUM_GROUPS=32. To prevent excessive index build times on AlleyTPI ($>12$ hours) for eu2005, dblp, and patents, we instead used MAX_L=4. Then for we decreased NUM_GROUPS to 16 for eu2005 and dblp and 8 for patents. We adjusted NUM_GROUPS more than MAX_L because small adjustments to the maximum pattern length greatly decrease the domain of patterns stored by the index (Kim et al., 2021).
@@
Experimental Setup. To reduce noise in our latency results, we repeat all inference results 3 times and report the median inference time. We do not do this for our cardinality estimates because this would unfairly reduce the impact of our sampling approach. These experiments are run on a server with an Intel(R) Xeon(R) CPU E7-4890 v2 @ 2.80GHz CPU, and all summary building and inference is done using a single thread. The reference implementation is available at: https://anonymous.4open.science/r/Cardinality-with-Colors-4333
@@ 8.1. Estimator Failure
There are two main ways that an estimator can fail to provide meaningful results: 1) it can time out which we define as taking longer than 1 minute to report a result 2) a sampling-based method can fail to find any qualifying samples. In Table 2, we show the proportion of queries that result in estimation failure for each dataset and technique. The simpler sampling-based methods (CS, WJ, JSUB, IMPR) face estimation failure even on the smaller, less dense query workloads (human, aids, lubm80) and fail to find samples for nearly any queries for the larger more complex workloads (yeast, dblp, youtube, eu2005, patents). Alley achieves much higher success rates across datasets due to its sophisticated sampling approach, but it still fails a significant portion of the time on four datasets. The summary-based methods (BSK++ and SumRDF), on the other hand, time out on over half of queries for four datasets. In contrast, because our approach applies sampling to a highly dense lifted graph, we never experience sampling failure on any query, across all workloads.
@@ 8.2. Accuracy
In Fig. 3, we show the relative error of various methods and workloads [^3] In these graphs, outliers (¿2 std. deviations) are shown as points. The inner box shows quartiles, and the whiskers are the max/min non-outlier values. Further, sample failure is treated as an estimate of 1.. Across workloads, our method, AvgMix32, is unbiased and scales well to larger, more cyclic workloads (yeast, dblp, youtube, eu2005, patents). It even achieves a median error of less than 2 on human, aids, dblp, and eu2005. We also reproduce the high accuracy of WanderJoin and Alley on the G-Care datasets. However, we find that all methods from (Park et al., 2020a) fail to scale to the larger more cyclic workloads. In particular, we reproduce the finding in (Kim et al., 2021) that WanderJoin, IMPR, and JSUB overwhelmingly fail to find a positive sample in a reasonable time on these datasets. Further, SumRDF times out on all larger queries due to its lack of aggregation and sampling.
@@
In this figure, we also vary the number of colors that we use for the lifted graph. Interestingly, increasing the number of colors kept does not straightforwardly improve accuracy. This is because we always use 500 samples during inference, so, as the lifted graph gets larger, the sampling procedure has a larger impact on the accuracy. Given this, we find that without increasing the sampling rate the optimal coloring uses either 32 or 64 colors.

### v2

@@ Abstract.
The first two authors contributed equally to this work.
PVLDB Reference Format:
PVLDB, 18(2): 130 - 143, 2024.
doi:10.14778/3705829.3705834 This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 18, No. 2 ISSN 2150-8097.
doi:10.14778/3705829.3705834
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣set%****␣main.tex␣Line␣25␣****https://github.com/uwdb/color.
@@ 4.2. Special Case: Stable Colorings
This theorem states that stable colorings are a perfect statistic for cardinality estimation of acyclic query graphs. However, we cannot use them in practice, because the number of stable colors needed to represent real graphs is close to the number of vertices in the data graph (Morris et al., 2021; Kayali and Suciu, 2022). Fortunately, we can prove an additional corollary for quasi-stable colorings:
Corollary 1.
Let $\epsilon$ be the approximation error of the lifted graph $\mathcal{G}$ defined as,
(6) $\epsilon=\max_{c_{1},c_{2}\in C}\frac{\tau_{\max}(c_{1},c_{2})}{\tau_{\min}(c_{1},c_{2})}$
Then, we can bound the error of our subgraph estimator as,
(7) $max(\frac{\Phi(Q,\mathcal{G})}{|hom(Q,G)|},\frac{|hom(Q,G)|}{\Phi(Q,\mathcal{G})})\leq\epsilon^{(|V_{Q}|-1)}$
Intuitively, as colorings approach stability, the estimate converges to the true subgraph count.
@@ 7.1. Partial Aggregation
Naively, the runtime of inference on the lifted graph is exponential in the size of the query graph, making it intractable for even moderately sized query graphs. The lifted graph is dense so the size of $Hom(Q,F)$ is approximately $|C|^{|Q|}$. Fortunately, we can rephrase Eq. 3 to drastically reduce this runtime via aggregate push-down. To do this, we express the set of lifted graph matches, $\text{hom}(Q,F)$, as $(c_{1},\ldots,c_{|Q|})\in C^{|Q|}$. We then use the definition of $W$ from 8 and, as before, we assume a topological ordering on the edges, $e_{1},\ldots,e_{|E|}$, and vertices, $v_{1},\ldots,v_{|Q|}$.
@@
At this point, we can identify this as a Functional Aggregate Query (Khamis et al., 2016) and apply the techniques there to solve it efficiently. [^2] Note, this is closely related to the variable elimination algorithm for probabilistic graphical models as well as tensor contraction algorithms. As an example, suppose that the query graph $Q$ is a line graph $v_{1}\rightarrow v_{2}\rightarrow v_{3}$. The naive expression would be the following $O(|C|^{3})$ expression,
@@
More generally, we can apply this strategy by choosing a variable order and at each step summing out the next variable in the order. The efficacy depends on the maximum number of variables present in any intermediate product. If an intermediate product involves $k$ variables, then we need to compute a relation of size $|C|^{k}$ which dominates the runtime. We defer the details of the proof to the technical report (tec, 2024), but this intuition can be formalized as follows using the theory of tree decompositions and treewidth,
@@
human 0.67 0.00 0.22 0.63 0.00 0.00 0.00 0.00 0.00 0.00 0.00
aids 0.69 0.07 0.14 0.28 0.00 0.04 0.01 0.00 0.02 0.39 0.00
lubm80 0.83 0.17 0.67 0.67 0.00 0.00 0.00 0.00 0.00 0.00 0.00
yeast 1.00 0.97 0.97 0.11 0.00 0.63 0.60 0.00 0.63 0.88 0.00
dblp 1.00 0.99 0.94 0.15 0.00 0.14 0.14 0.00 0.70 0.85 0.00
youtube 0.99 0.93 1.00 0.22 0.00 0.10 0.05 0.00 0.63 0.78 0.00
eu2005 0.95 0.90 0.91 0.55 0.00 0.00 0.00 0.00 0.22 0.44 0.00
patents 0.98 0.88 0.98 0.08 0.00 0.13 0.13 NA 0.67 0.79 0.00
Of course, this relies on finding a good ordering of the query vertices, and finding the optimal one is naively NP-Hard with respect to the size of $Q$. Fortunately, there are very effective heuristics for identifying good tree decompositions, and we apply these in our system, using the min-fill heuristic (Rollon and Larrosa, 2011). To accommodate our sampling techniques, we restrict these tree decompositions to path decompositions and get results relative to the pathwidth.
7.2. Sampling Techniques
In real systems, cardinality estimation needs to be extremely fast and consistent because its overhead is seen by every query. While partial aggregation speeds up inference for simple query graphs with low treewidth, larger and denser query graphs still pose a problem. To avoid this, we use a sampling procedure that ensures linear inference w.r.t query size (see Fig. 13). This is similar to a weighted version of the WanderJoin algorithm from (Li et al., 2019). We apply a Thompson-Horowitz estimator to random paths within the lifted graph. However, we adjust the method in two important ways: 1) we incorporate the sampling into the aggregation framework from Sec. 7.1 2) we apply importance sampling to account for the fact that different paths within the lifted graphs contribute more or less to the cardinality estimate.
Sampling During Aggregation
At each step, we process a vertex of the query graph and materialize an intermediate result consisting of partial colorings and their weights. After materialization, we apply sampling in order to reduce the amount of partial colorings that we extend in the next step. By doing this at each step, we can maintain a constant number of partial colorings at all times and ensure a linear runtime w.r.t. the size of the query graph. Because we still apply aggregation, we reduce the amount of sampling required.
Importance Sampling
This is a classical technique for approximating the value of an integral, and we adapt it here by noting that our summation in 10 is a discrete integral over a product. The core idea is to sample points of the integrand which contribute more heavily to the result with higher probability in order to reduce the variance. Because determining the contribution of a partial coloring to the final result is challenging, we approximate this contribution via its partial count. This assumes that a high partial count leads to a high contribution to the final sum. To keep our estimator unbiased, we apply Thompson-Horowitz estimation and scale each sampled partial count by the inverse of its selection probability. Finally, we scale the total sampled weight to set it to the total weight prior to sampling.
@@ 7.3. Handling Updates
Updates pose a challenge to summary-based estimators because the statistics which they collect become stale over time as updates are applied to the database. Traditional estimators simply rebuild the summary intermittently(Haas, 1999). This leads to severe decreases in accuracy under even modest updates because the estimator is entirely ”blind” to them. On the other hand, recalculating the summary before each query is very costly. In this section, we demonstrate how COLOR supports a middle ground approach that applies fast, basic updates to the lifted graph, allowing it to maximize the time between full rebuilds.
@@
Given an update $\theta_{E}=(v_{1},v_{2},A)$, the edge update function $\delta_{E}$ marginally increases $\tau$ for the combination of attributes and colors in the edge update. However, the edge update doesn’t directly contain the colors of $v_{1}$ and $v_{2}$ or the attributes of $v_{2}$. To retrieve the colors, $c_{1}$ and $c_{2}$, associated with $v_{1}$ and $v_{2}$, we look up their values in $\pi$ which we store compactly (and approximately) as a series of cuckoo filters. To calculate the attributes of $v_{2}$, we leverage statistics about the attribute distribution in $c_{2}$ to sample a set of attributes.
@@ 8. Evaluation
    Never experiences estimation failure on any query across all workloads.
@@
    Produced up to $\mathbf{10^{7}\times}$ tighter bounds when using $\tau_{max}$ than BoundSketch while being many orders of magnitude faster.
@@
Datasets & Workload. We consider datasets from (Park et al., 2020a) and (Sun and Luo, 2020) for our analysis. These datasets come from a variety of domains. Those from (Sun and Luo, 2020) are undirected graphs whose queries are larger and vary in density. Those from (Park et al., 2020a) are directed graphs whose queries are smaller and vary in shape. We adapt undirected workloads to directed methods by including reverse edges in the data graph but not in the query graphs. Both of these sources include label predicates in their queries with the former using both edge and vertex labels and the latter using only vertex labels. Table 3 shows their different characteristics where $|\ell_{V}|$ and $|\ell_{E}|$ are the number of edge and vertex labels.
Comparison Methods. For comparison, we use a superset of the methods considered in (Park et al., 2020a) and additionally apply them to the larger, more complex datasets from (Sun and Luo, 2020). These methods include: 1) Correlated Sampling (CS) (Vengerov et al., 2015) 2) Characteristic Sets (CSet)(Neumann and Moerkotte, 2011) 3) Wander Join (WJ) (Li et al., 2019) 4) Alley (alley) and alleyTPI (Kim et al., 2021) 5) Join Sampling with Upper Bounds (JSUB), an adaptation of (Zhao et al., 2018) 6) Bound Sketch (BSK) (Cai et al., 2019) which corresponds to a hash-based coloring and using the max degree estimator. We further apply our partial aggregation optimization, and we call this improved version BSK++. 7) IMPR (Chen and Lui, 2018) and 8) SumRDF (Stefanoni et al., 2018) 9) Learned Sketch for Subgraph Counting (LSS) (Zhao et al., 2023), a query-driven deep learning approach. Due to hardware compatibility issues, this method was run on a different machine than the others with an Intel(R) Core(TM) i7-9750H CPU @ 2.60GHz CPU. This machine had 32GB of memory, resulting in an OOM error for the patents dataset which we treat as NA rather than estimator failure. 10) We also include a traditional independence-based estimator (IndEst) corresponding to Def. 2.4. For the sampling based estimators, we apply the default sampling ratios from (Park et al., 2020a) and (Kim et al., 2021) (i.e. .03 for all methods except for Alley which uses .001). AlleyTPI uses a maximum pattern length (MAX_L) and a maximum number of stored label groups (NUM_GROUPS) when building its index, with default values of MAX_L=4 or MAX_L=5 depending on the dataset, and NUM_GROUPS=32. To prevent excessive index build times on AlleyTPI ($>12$ hours) for eu2005, dblp, and patents, we used MAX_L=4. We decreased NUM_GROUPS to 16 for eu2005 and dblp and 8 for patents. We adjusted NUM_GROUPS more than MAX_L because small adjustments to the maximum pattern length greatly decrease the domain of patterns stored by the index (Kim et al., 2021). As in the original work, we get estimates for LSS via a 5-fold cross-validation where 4/5ths of the queries are trained on and 1/5 is estimated during each fold. Lastly, recently, there has been work on the isomorphism variant of subgraph cardinality estimation from both the deep learning and sampling perspective(Shin et al., 2024; Wang et al., 2022). Unfortunately, this difference in problem setting makes them incomparable with the other methods considered here, so we exclude them from our evaluation.
@@
Experimental Setup. To reduce noise, we repeat all inference results 3 times and report the median inference time. We do not do this for our cardinality estimates because this would unfairly reduce the impact of our sampling approach. These experiments are run on a server with an Intel(R) Xeon(R) CPU E7-4890 v2 @ 2.80GHz CPU, and all summary building and inference is done using a single thread, unless stated otherwise. The reference implementation is available at: https://github.com/uwdb/color
@@ 8.1. Estimator Failure
There are two ways that an estimator can fail to provide meaningful results: 1) time outs, which we define as taking longer than 1 minute to report a result 2) sampling failure, not finding any qualifying samples. In Table 2, we show the proportion of queries that result in estimation failure for each dataset and technique. Simpler sampling-based methods (CS, WJ, JSUB, IMPR) face estimation failure even on the smaller query workloads and fail to find any samples most queries on the larger workloads. Alley achieves much higher success rates due to its sophisticated sampling approach but still fails a significant portion of the time on four datasets. Summary-based methods (BSK++ and SumRDF), on the other hand, time out on over half of queries for four datasets. In contrast, because our approach applies sampling to a highly dense lifted graph, we never experience sampling failure on any query, across all workloads. As the data graph contains hundreds of thousands of edges or more, which are mapped into a lifted graph with at most 32 nodes (colors), the probability that any two colors have an edge connecting them is high.
@@ 8.2. Accuracy
In Fig. 3, we show the relative error of various methods and workloads [^3] In these graphs, outliers (¿2 std. deviations) are shown as points. The inner box shows quartiles, and the whiskers are the max/min non-outlier values. Further, sample failure is treated as an estimate of 1.. Relative error is defined as the ratio of the estimate to the true cardinality. Scores greater than one indicate an overestimate, those under an underestimate. Across workloads, our method, AvgMix32, is unbiased and scales well to larger, more cyclic workloads (yeast, dblp, youtube, eu2005, patents). It even achieves a median error of less than 2 on human, aids, dblp, and eu2005. We also reproduce the high accuracy of WanderJoin and Alley on the G-Care datasets. However, we find that all methods from (Park et al., 2020a) fail to scale to the larger more cyclic workloads. In particular, we reproduce the finding in (Kim et al., 2021) that WanderJoin, IMPR, and JSUB overwhelmingly fail to find a positive sample in a reasonable time on these datasets, and that LSS achieves reasonable accuracy on a variety of workloads. Further, SumRDF times out on all larger queries.
@@
In this figure, we also vary the number of colors that we use for the lifted graph. Interestingly, increasing the number of colors kept does not straightforwardly improve accuracy. This is because we always use 500 samples during inference. As the lifted graph gets larger, the sampling procedure has a larger impact. Given this, we find that the optimal coloring uses either 32 or 64 colors.